Reports 1-1 of 1 Clear search Modify search
MIF (General)
takafumi.ushiba - 9:44 Wednesday 27 November 2019 (11783) Print this report
BS TML damping filter OFF during FPMI

[Takano, Ushiba]

We found that BS_TM_DAMP_L was engaged when the FPMI was locked.
This filter could be induced control noise at MICH signals and it is not necessary during MICH control.
So, we turned off the filter and see the spectra (Attached).

We are not sure it affected the sensitivity but today's morining sensitivity is smaller than that on November 14 around 10 Hz and 20 - 40 Hz.
I have already implemented turning off the BS_TM_DAMP_L filter  at TRANSITION_TO_REFL17 and turning on the filter at DOWN state of the guardian.

Images attached to this report
Search Help